The Bouchard Company's EPS was $5.39 in 2016, up from $3.08 in 2011. The company pays out 55% of its earnings as dividends, and its common stock sells for $40. Calculate the past growth rate in earnings. (Hint: This is a 5-year growth period.) Round your answer to two decimal places. % The last dividend was D0 = 0.55($5.39) = $2.96. Calculate the next expected dividend, D1, assuming that the past growth rate continues. Do not round off intermediate calculations. Round your answer to the nearest cent. $ What is Bouchard's cost of retained earnings, rs? Round your answer to two decimal places. Do not round your intermediate calculations. %
Answer #1
Annual average growth rate
=((last value/First value)^(1/Time between 1st and last value)-1)*100
=((5.39/3.08)^(1/5)-1)*100
Annual Growth rate% = 11.84

Next dividend = EPS*payout ratio*(1+growth rate)

=5.39*0.55*(1+0.1184)=3.32

As per DDM
Price = Dividend in 1 year/(cost of equity - growth rate)
40 = 3.3154968/ (Cost of equity - 0.1184)
Cost of equity% = 20.13